# Scaling of Disorder Operator and Entanglement Entropy at Easy-Plane Deconfined Quantum Criticalities

### v1

@@ III Disorder Operator
where all the coefficients are functions of $\theta$ and the log-coefficient $s$ follows a universal function of both $\theta$ and the opening angles of the corners of $M$ ($\pi/2$ for rectangle region). Given a smooth region (without a corner), there should be no corner correction and the log-coefficient $s=0$. Such area-law decay of disorder operator, i.e. $\ln|X_{M}(\theta)|=-a_{1}L+a_{0}$, holds both at the QCP and inside the gapped symmetric phases.
To detect the scaling of the disorder operator at the quantum phase transition point of EPJQ model in Eq. (1), we choose the entanglement region $M$ to be a $L\times L/2$ cylinder region (without corner) in the lattice, as shown in the right inset of Fig. 1 (a), with system size for $L=8$ up to $L=144$. For a good comparison, we calculated the U(1) disorder operator (with small rotation angle $\theta=0.2$) for the Néel phase (standard spin-1/2 Heisenberg model on the square lattice), the O(3) QCP of $J_{1}-J_{2}$ model in Eq. (6), DQCP of JQ3 model (Eq. (1) at $\Delta=1$), and quantum phase transition points of EPJQ of Eq. (1) for $\Delta=0,0.1,0.3,0.5$, respectively. The results are shown in Fig. 3 (a).
@@ IV Entanglement Entropy
Next we investigate the EE at the transition points of various $\Delta$ of Eq. (1). To this end, we employ the nonequilibrium increment method Alba (2017); D’Emidio (2020); Zhao et al. (2022a, b) within the framework of SSE-QMC simulation Sandvik (1999); Syljuåsen and Sandvik (2002) to determine the second Rényi entropy of the EPJQ model at phase transition points for various values of $\Delta$. Rényi entanglement entropy is defined as $S_{A}^{(n)}=\frac{1}{1-n}\ln\operatorname{Tr}\rho_{A}^{n}$ which can be reexpressed in the form of $S_{A}^{(n)}=\frac{1}{1-n}\ln\frac{Z_{A}^{(n)}}{Z^{(n)}}$ according to the replica trick Calabrese and Cardy (2004). The nonequilibrium method is based on Jarzynski’s equality Jarzynski (1997), which relates the free energy difference between two systems with the total work done during a tuning process from one system to another. We regard the partition functions $Z_{A}^{(n)}$ and $Z^{(n)}$ as those of two different physical systems, then it is natural to apply the Jarzynski’s equality and design a tunning process between the two systems to calculate the Rényi entropies Alba (2017); D’Emidio (2020). In practice, we follow the incremental version Zhao et al. (2022a, b) of Ref. D’Emidio (2020), and it can overcome the fact that the EE is in general an exponential observable Zhang et al. (2024); Zhou et al. (2024). We conduct the simulation on a $L\times L$ square lattice with periodic boundary conditions, and fix with $\beta=L$ and choose the subregion A to be a $L\times L/2$ cylinder defined same as region $M$ in the right inset of Fig. 1 (a), which has no sharp corners on the entanglement boundary.
For $\Delta<1.0$, in the AFXY phase, the system spontaneously breaks the SO(2) spin symmetry and exhibits one Goldstone mode. In this case, the entanglement entropy (EE) is expected to scale as Metlitski and Grover (2015)
@@
where $N_{g}$ represents the number of Goldstone modes. Eq. (9) has been verified for both XY and Néel ordered states, which have one and two Goldstone modes Helmes and Wessel (2014); Kulchytskyy et al. (2015); D’Emidio (2020); Zhao et al. (2022b); Song et al. (2023a); Deng et al. (2023b); Zhou et al. (2024); D’Emidio et al. (2024), respectively. We first test our algorithm in the AFXY phase of the EPJQ model at $\Delta=0$ and $q=0.3\ll q_{c}$, with the results displayed in Fig. 4. We measure the second order Rényi EE for the subsystem $M$. The system sizes utilized are $L=8,12,16,\ldots,32$ and the temperature is fixed at $\beta=L$. As shown in Fig. 4(b), by fitting to Eq. (9), we observe a good agreement between our data with the scaling form of Eq. (9). Moreover, the log-coefficient obtained from curve fitting is $\frac{N_{G}}{2}=0.497\pm 0.011$, aligning well with the expectation of existence of one Goldstone mode, which corresponds to spontaneous SO(2) symmetry breaking. The analysis of the fitting quality can be found in the Sec. II in SM sup .
We then proceed to the phase transition points. Our extrapolated finite order parameter data in Fig. 1(b) shows that the first-order behavior is more pronounced as $\Delta$ decreases from 1 to 0. However, the relation between the scaling of entanglement entropy (EE) with the strengthen of first-order behavior remains unexplored. We measure the second Rényi EE at the transition points in Table. 1 for various values of $\Delta$ and system sizes $L=8,12,16,\ldots,44$ and fix $\beta=L$. The results are presented in Fig. 5. For all examined values of $\Delta$, we observe a clear logarithmic correction to the area law term in the scaling of EE (Eq. (9)). Additionally, as $\Delta$ decreases and the first-order behavior strengthens, the fitted log-coefficient $b$ increases. Interestingly, for $\Delta=0.1$ and $\Delta=0.3$, the fitted value of $b$ is even consistent with 0.5, which corresponds to the existence of a single Goldstone mode, the same as the scaling of EE deep inside the AFXY phase in Fig. 4. Our findings suggest that at least for $\Delta\leq 0.3$ the log-coefficient $b$ characterizes a evident first order transition with SO(2) symmetry breaking associated with one Goldstone mode.
At $\Delta=1.0$, where the system reduces to the isotropic JQ3 model, it has been discovered that even with a smooth entanglement boundary, the system exhibits anomalous positive logarithmic subleading corrections, resembling the Goldstone mode contribution, while the coefficient is only around 0.2-0.3 Song et al. (2023a, b). In addition, a recent study Deng et al. (2024) considers carefully the finite-size effects in Eq. (9), and use the modified fitting function for Goldstone mode phase
@@
where $\rho_{s}(L)$ is the finite-size spin stiffness and $I(L)$ the finite-size transverse spin susceptibility. Through this procedure, the log-coefficient is fitted to be around 2 which corresponds to four Goldstone modes Deng et al. (2024).This is used to suggest that the system breaks the emergent SO(5) symmetry at the transition point, and this process can be well detected by the scaling of EE even at relatively small system sizes.
We followed this procedure, calculated the finite-size spin stiffness $\rho_{s}(L)$ and transverse spin susceptibility $I(L)$, and substituted them in Eq. (10), same as in Ref. Deng et al. (2024) to fit our EE data. However, as shown in Fig. 6, we find it is not so easy to observe the same convergence of log-coefficient to 2 at $\Delta=1.0$ or to 1.5 at $\Delta<1.0$, which would suggest the breaking of emergent SO(5) (four Goldstone modes) and SO(4) (three Goldstone modes) symmetries, respectively. In fact, at $\Delta=1.0$ the last two data points with larger $L_{\text{min}}$ are quite close to $b=2$, although the last data point slightly deviates below $b=2$.
For $\Delta<1.0$, except for $\Delta=0.3$, the last data points all seem to be consistent with $b=1.5$. However, within our limited data quality, we do not observe the stability of fitted results with respect to $L_{\text{min}}$ for any $\Delta$. It is possible that our results can be affected by stronger finite-size effects as we fixed $\beta=L$ to approach the thermodynamic limit, while Ref. Deng et al. (2024) used the projector quantum Monte Carlo Sandvik (2010, 2005) to simulate the ground state of the system and fix the projection power to be $m=L^{3}$. But we think physically the more important point is, that we observed at $\Delta<1.0$ the first order is enhanced, so that the emergent $SO(4)$ behavior can be weakened at the system sizes we considered, and a strong first-order behavior with only one Goldstone mode might be mixed with the $SO(4)$ symmetry breaking behavior with three Goldstone modes. In this case, the fitted log coefficient with Eq. (10) for $\Delta<1.0$ could drift significantly from 1.5 to 0.5 when larger system sizes are considered. We therefore think that overall our fitting results with Eq. (10) are consistent with Ref. Deng et al. (2024) for JQ3 model and show the capability of EE in characterizing the emergent symmetry and weakly first order behaviors at the complicated DQCPs.
@@ V Discussion
Our work shows the two non-local measurements – the disorder operator and entanglement entropy – consistently exhibit anomalous scaling behaviors at the easy-plane deconfined quantum criticalities. By adjusting the anisotropy from the Heisenberg limit ($\Delta=1$) to the easy-plane limit ($\Delta=0$), we find that as the first-order nature of the transition is being amplified, the anomalous behavior of both the disorder operator and entanglement entropy becomes more apparent, resembling contributions from Goldstone modes. Interestingly, when $\Delta\leq 0.3$, the log coefficient in EE seem to converge to $b=0.5$ which is consistent with one Goldstone mode and SO(2) symmetry breaking at the first-order transition points. Additionally, when applying the finite-size fitting form of Goldstone mode phase (Eq. (10)) to fit our EE data, we obtain the enhanced log coefficients close to that of emergent SO(5) symmetry breaking at $\Delta=1$ and emergent SO(4) symmetry breaking at $\Delta<1$. However, the log coefficients do not show the good convergence behavior, as shown in Ref. Deng et al. (2024), which maybe explained by insufficient data quality or very weak emergent SO(4) symmetry at $\Delta<1$ due to the strengthening of first order behavior. Our work thus provides strong evidence that the observed anomalous scaling behaviors of the entanglement measurements (disorder operator and EE) indeed come from the weakly-first-order nature of DQCPs of JQ model realizations both at $\Delta=1$ and $\Delta<1$.
@@
Note that our work and previous studies Wang et al. (2022); Zhao et al. (2022a); Song et al. (2023a, b); Liu et al. (2023); Liao et al. (2023); Liu et al. (2024); D’Emidio and Sandvik (2024); Deng et al. (2024) on the EE scaling of smooth cut on a square lattice are not contradicted with the recent work by D’Emidio and Sandvik D’Emidio and Sandvik (2024) where they consider a $\frac{\pi}{4}$-degrees-tilted square lattice and observe the normal scaling of EE obeying the requirements of CFTs. We think the cut dependence of EE actually supports the transition is indeed weakly-first order and the tilted cut and standard cut possibly just capture the two sides of the JQ-type DQCP story. On the one hand, it is a weakly first transition which can be detected by the standard cut, and on the other hand the system might be close to a real continuous transition so that the tilted cut captures some essential information of the CFT at small system sizes. In conclusion, scaling of these nonlocal observables have proven to be both powerful and sensitive tools in characterising this kind of weakly first-order transitions Wang et al. (2022); Zhao et al. (2022a); Song et al. (2023a); Deng et al. (2024), and it provides strong evidence that the JQ type DQCP is weakly first order however very close to a real continuous transition Takahashi et al. (2024), possibly of multicritical type related with the SO(5) model with Wess-Zumino-Witten term Zhao et al. (2020); Chen et al. (2023); Chester and Su (2024); Takahashi et al. (2024); Chen et al. (2024) or the ‘walking" pseudo-criticality behavior at the transition Wang et al. (2017); Nahum (2020); Ma and Wang (2020); Zhou et al. (2023).

### v2

@@ III Disorder Operator
where all the coefficients are functions of $\theta$ and the log-coefficient $s$ follows a universal function of both $\theta$ and the opening angles $\phi$ of the corners of $M$ ($\phi=\pi/2$ for rectangle region). Given a smooth region M (without corners on boundary), there should be no corner correction and the log-coefficient $s=0$. Such area-law decay of disorder operator, i.e. $\ln|X_{M}(\theta)|=-a_{1}L+a_{0}$, holds both at the QCP and inside the gapped symmetric phases.
To detect the scaling of the disorder operator at the quantum phase transition point of EPJQ model in Eq. (1), we choose the entanglement region $M$ to be a $L\times L/2$ cylinder region (without corners) in the lattice, as shown in the right inset of Fig. 1 (a), with system size for $L=8$ up to $L=144$. For a good comparison between different quantum states, we calculated the U(1) disorder operator (with small rotation angle $\theta=0.2$) for the Néel phase (standard spin-1/2 Heisenberg model on the square lattice), the O(3) QCP of $J_{1}-J_{2}$ model in Eq. (6), DQCP of JQ3 model (Eq. (1) at $\Delta=1$), and quantum phase transition points of EPJQ of Eq. (1) for $\Delta=0,0.1,0.3,0.5$, respectively. The results are shown in Fig. 3 (a).
@@ IV Entanglement Entropy
Next we investigate the EE at the transition points of various $\Delta$ of Eq. (1). To this end, we employ the nonequilibrium increment method Alba (2017); D’Emidio (2020); Zhao et al. (2022a, b) within the framework of SSE-QMC simulation Sandvik (1999); Syljuåsen and Sandvik (2002) to determine the second Rényi entropy of the EPJQ model at phase transition points for various values of $\Delta$. Rényi entanglement entropy is defined as $S_{A}^{(n)}=\frac{1}{1-n}\ln\operatorname{Tr}\rho_{A}^{n}$ which can be reexpressed in the form of $S_{A}^{(n)}=\frac{1}{1-n}\ln\frac{Z_{A}^{(n)}}{Z^{(n)}}$ according to the replica trick Calabrese and Cardy (2004). The nonequilibrium method is based on Jarzynski’s equality Jarzynski (1997), which relates the free energy difference between two systems with the total work done during a tuning process from one system to another. We regard the partition functions $Z_{A}^{(n)}$ and $Z^{(n)}$ as those of two different physical systems, then it is natural to apply the Jarzynski’s equality and design a tunning process between the two systems to calculate the Rényi entropies Alba (2017); D’Emidio (2020). In practice, we follow the incremental version Zhao et al. (2022a, b) of Ref. D’Emidio (2020), and it can overcome the obstacles that the EE is in general an exponential observable Zhang et al. (2024); Zhou et al. (2024). We conduct the simulation on a $L\times L$ square lattice with periodic boundary conditions, and fix with $\beta=L$ and choose the subregion A to be a $L\times L/2$ cylinder defined same as region $M$ in the right inset of Fig. 1 (a), which has no sharp corners on the entanglement boundary.
For $\Delta<1.0$, in the AFXY phase, the system spontaneously breaks the SO(2) spin symmetry and possesses one Goldstone mode. In this case, EE is expected to scale as Metlitski and Grover (2015)
@@
where $N_{g}$ represents the number of Goldstone modes. Eq. (9) has been extensively verified for both XY phase with one Goldstone mode and Néel ordered states with two Goldstone modes respectively Helmes and Wessel (2014); Kulchytskyy et al. (2015); D’Emidio (2020); Zhao et al. (2022b); Song et al. (2023a); Deng et al. (2023b); Zhou et al. (2024); D’Emidio et al. (2024) . We first test our algorithm in the AFXY phase of the EPJQ model at $\Delta=0$ and $q=0.3$ deep in the AFMXY phase, with the results displayed in Fig. 4. We measure the second order Rényi EE for the $L\times L/2$ subsystem $A$ on a $L\times L$ square lattice with smooth boundary. The system sizes utilized are $L=8,12,16,\ldots,32$ and the temperature is fixed at $\beta=L$. As shown in Fig. 4(b), by fitting the finite-size EE data to Eq. (9), we observe a good agreement between our data with the scaling form of Eq. (9). Moreover, the log-coefficient obtained from curve fitting is $\frac{N_{G}}{2}=0.497\pm 0.011$, aligning well with the expectation of existence of one Goldstone mode, which corresponds to spontaneous SO(2) symmetry breaking. The analysis of the fitting quality can be found in the Sec. II in SM sup .
We then proceed to the phase transition points. Our extrapolated finite order parameter data in Fig. 1(b) shows that the first-order behavior is more pronounced as $\Delta$ decreases from 1 to 0. However, the relation between the scaling of entanglement entropy (EE) with the strengthen of first-order behavior remains unexplored. We measure the second Rényi EE at the transition points in Table. 1 for various values of $\Delta$ and system sizes $L=8,12,16,\ldots,44$ and fix $\beta=L$. The subsystem A is again chosen to be $L\times L/2$ cylinder withour corners on the boundary. As presented in Fig. 5, for all examined values of $\Delta$, we observe a clear logarithmic correction to the area law term in the scaling of EE (Eq. (9)). Additionally, as $\Delta$ decreases and the first-order behavior strengthens, the fitted log-coefficient $b$ increases. Interestingly, for $\Delta=0.1$ and $\Delta=0.3$, the fitted value of $b$ is even consistent with 0.5, which corresponds to the existence of a single Goldstone mode, the same as the scaling of EE deep inside the AFXY phase in Fig. 4. Our findings suggest that at least for $\Delta\leq 0.3$ the log-coefficient $b$ characterizes a evident first order transition with SO(2) symmetry breaking associated with one Goldstone mode.
At $\Delta=1.0$, where the system reduces to the isotropic JQ3 model, it has been discovered that even with a smooth entanglement boundary, the system exhibits anomalous positive logarithmic subleading corrections, resembling the Goldstone mode contribution, while the coefficient is only around 0.2-0.3 Song et al. (2023a, b). In addition, a recent study Deng et al. (2024) considers carefully the finite-size effects in scaling form for Goldstone mode phase, and apply the modified fitting function
@@
where $\rho_{s}(L)$ is the finite-size spin stiffness and $I(L)$ the finite-size transverse spin susceptibility which are measured separately in QMC simulations and substituted in the above relation to fit the finite-size EE data. Through this procedure, the log-coefficient is fitted to be around 2 which corresponds to four Goldstone modes Deng et al. (2024). This is used to suggest that the system breaks the emergent SO(5) symmetry at the transition point, and this process can be well detected by the scaling of EE even at relatively small system sizes Deng et al. (2024).
We followed this procedure, calculated the finite-size spin stiffness $\rho_{s}(L)$ and transverse spin susceptibility $I(L)$, and substituted them in Eq. (10), same as in Ref. Deng et al. (2024) to fit our EE data. However, as shown in Fig. 6, we find it is difficult to observe the same convergence of log-coefficient to 2 at $\Delta=1.0$ or to 1.5 at $\Delta<1.0$, which would suggest the breaking of emergent SO(5) (four Goldstone modes) and SO(4) (three Goldstone modes) symmetries, respectively. In fact, at $\Delta=1.0$ the last two data points with larger $L_{\text{min}}$ are quite close to $b=2$, although the last data point slightly deviates below $b=2$.
For $\Delta<1.0$, except for $\Delta=0.3$, the last data points all seem to be consistent with $b=1.5$. However, within our limited data quality, we do not observe the stability of fitted results with respect to $L_{\text{min}}$ for any $\Delta$. It is possible that our results can be affected by stronger finite-size effects as we fixed $\beta=L$ to approach the thermodynamic limit, while Ref. Deng et al. (2024) used the projector quantum Monte Carlo Sandvik (2010, 2005) to simulate the ground state of the system and fix the projection power to be $m=L^{3}$. Another possible explanation is, as we observed at $\Delta<1.0$ the first order is enhanced, the emergent $SO(4)$ behavior can be weakened and broken into remaining VBS order and Goldstone phase at small system sizes. In this case, within the system sizes we considered, an evident first-order behavior with only one Goldstone mode might be mixed with the $SO(4)$ symmetry breaking behavior with three Goldstone modes. Then the fitted log coefficient with Eq. (10) for $\Delta<1.0$ could drift significantly from 1.5 to 0.5 as system sizes grow. Therefore our fitting results with Eq. (10) are consistent with Ref. Deng et al. (2024) for JQ3 model and show the capability of EE in characterizing the emergent symmetry and weakly first order behaviors at the complicated DQCPs.
@@ V Discussion
Our work shows the two non-local measurements – the disorder operator and entanglement entropy – consistently exhibit anomalous scaling behaviors at the easy-plane deconfined quantum criticalities. By adjusting the anisotropy from the Heisenberg limit ($\Delta=1$) to the easy-plane limit ($\Delta=0$), we find that as the first-order nature of the transition is amplified, the anomalous behavior for both disorder operator and entanglement entropy becomes more apparent, resembling contributions from Goldstone modes. Interestingly, when $\Delta\leq 0.3$, the log coefficient in EE seem to converge to $b=0.5$ which is consistent with one Goldstone mode and SO(2) symmetry breaking at the first-order transition points. Additionally, when applying the finite-size fitting form of Goldstone mode phase (Eq. (10)) to fit our EE data, we obtain the enhanced log coefficients close to that of emergent SO(5) symmetry breaking at $\Delta=1$ and emergent SO(4) symmetry breaking at $\Delta<1$. However, the log coefficients do not show the good convergence behavior, as shown in Ref. Deng et al. (2024), which maybe explained by insufficient data quality or very weak emergent SO(4) symmetry at $\Delta<1$ due to the strengthening of first order behavior. Our work thus provides strong evidence that the observed anomalous scaling behaviors of the entanglement measurements (disorder operator and EE) indeed come from the weakly-first-order nature of DQCPs of JQ model realizations both at $\Delta=1$ and $\Delta<1$.
@@
Note that our work and previous studies Wang et al. (2022); Zhao et al. (2022a); Song et al. (2023a, b); Liu et al. (2023); Liao et al. (2023); Liu et al. (2024); D’Emidio and Sandvik (2024); Deng et al. (2024) on the EE scaling of smooth cut on a square lattice are not contradicted with the recent work by D’Emidio and Sandvik D’Emidio and Sandvik (2024) where they consider a $\frac{\pi}{4}$-degrees-tilted square lattice and observe the normal scaling of EE obeying the requirements of CFTs. On the contrary, the cut dependence of EE might support the transition is indeed weakly-first order and the tilted cut and standard cut possibly just capture the two sides of the JQ-type DQCP story. On the one hand, it is a weakly first transition which can be detected by the standard cut, and on the other hand the system might be close to a real continuous transition so that the tilted cut captures some essential information of the CFT at small system sizes. In conclusion, scaling of these nonlocal observables have proven to be both powerful and sensitive tools in characterising this kind of weakly first-order transitions Wang et al. (2022); Zhao et al. (2022a); Song et al. (2023a); Deng et al. (2024), and it provides strong evidence that the JQ type DQCP is weakly first order however very close to a real continuous transition Takahashi et al. (2024), possibly of multicritical type related with the SO(5) model with Wess-Zumino-Witten term Zhao et al. (2020); Chen et al. (2023); Chester and Su (2024); Takahashi et al. (2024); Chen et al. (2024) or the ‘walking" pseudo-criticality behavior at the transition Wang et al. (2017); Nahum (2020); Ma and Wang (2020); Zhou et al. (2023).